# Reinforcing Question Answering Agents with Minimalist Policy Gradient Optimization

### v2

@@
Abstract
Large Language Models (LLMs) show remarkable versatility but face challenges in Question Answering (QA) due to knowledge cutoffs and hallucination. While Retrieval-Augmented Generation (RAG) helps by integrating external knowledge, current methods often depend on in-context learning, which can be limited by backbone models and prompt engineering. To overcome these limitations, we introduce Mujica (Multi-hop Joint Intelligence for Complex Question Answering), a novel multi-agent collaborative framework employing a divide-and-conquer strategy for QA. This framework is trained using MyGO (Minimalist policy Gradient Optimization), a new reinforcement learning method that directly samples and learns from the optimal policy. Empirical results across multiple datasets demonstrate that the Mujica-MyGO approach significantly enhances multi-hop QA performance for various LLMs.
1 Introduction
@@
The advent of Large Language Models (LLMs) ouyang2022training ; achiam2023gpt has redefined expectations for Artificial Intelligence, offering new opportunities for Question Answering (QA) tasks due to their impressive language understanding and generation capabilities. However, directly applying LLMs to QA is challenging because of hallucination huang2025survey and knowledge cutoffs cheng2024dated . While injecting knowledge through continue learning is a possible solution, it is computationally expensive and risks catastrophic forgetting kirkpatrick2017overcoming , where models lose previously acquired knowledge. To address this issue, Retrieval-Augmented Generation (RAG) guu2020retrieval ; lewis2020retrieval , a technique enabling LLMs to incorporate external knowledge via prompts, has drawn significant attention among researchers.
Compared with single-hop questions which have one-turn retrieve-then-answer pipeline, most advanced RAG systems PoG ; Search-o1 ; RoG ; shao2023enhancing ; khattab2022demonstrate adopt a multi-turn pipeline to effectively address multi-hop questions. This involves decomposing the question, retrieving information for each sub-question, and synthesizing a final answer. Such a system can be naturally conceptualized as an agent taking charge of making a plan on how to answer the question, interacting with external data sources, and conducting necessary reasoning ToG ; ToG2 ; Search-o1 . However, existing approaches primarily rely on in-context demonstration as the sole guidance. However, in-context learning (ICL) ICL has several intrinsic limitations in the QA task. First, including few-shot demonstrations requires the LLM to process more tokens for every question, which not only reduces inference speed but also increases the risk of performance degradation due to attention dispersion caused by excessively sequence lengths. Second, for different types of complex questions, one may need to provide demonstrations tailored to the specific question type to ensure effectiveness. Furthermore, in-context learning operates as a form of analogy-based reasoning, which does not improve the underlying fundamental capability of LLMs to handle QA tasks. Consequently, the model may struggle with uncommon questions in specialized domains. Therefore, to develop a truly powerful and efficient QA agent, applying post-training on LLMs becomes essential.
Training QA agents remains a challenging task. While ground truth answers are typically available, the golden reasoning paths required to derive these answers are often absent in real-world scenarios and mainstream QA datasets yang2018hotpotqa ; trivedi2021musique ; QALD-10 . Due to the lack of intermediate supervision signals, directly applying supervised finetuning (SFT) on QA agents becomes infeasible. A naive and intuitive solution is to employ larger-scale LLMs to generate reasoning paths and transfer the knowledge into smaller models through knowledge distillation. However, this approach is computationally expensive and lacks scalability, making it impractical for further improving the most advanced LLMs. Considering the aforementioned limitations, reinforcement learning (RL) emerges as a promising alternative, allowing LLMs to autonomously improve their capabilities. Recent works RAG-Gym ; IM-RAG proposes to employ the intrinsic reasoning capability of LLMs to generate “pseudo” golden reasoning paths. As training progresses, the quality of the generated reasoning paths generally improves, gradually approaching the effectiveness of knowledge distillation.
Figure 1: The end-to-end architecture of the proposed Mujica framework.
Nevertheless, directly employ existing RL solutions for LLM post-training (e.g., PPO schulman2017proximal and GRPO shao2024deepseekmath ) on complex question answering tasks is resource-intensive and inefficient. In order to stabilize the training process, these approaches schulman2017proximal ; shao2024deepseekmath ; team2025kimi ; guo2025deepseek apply gradient rescaling for individual data samples, which is primarily achieved through importance sampling and clipping mechanism. However, gradient rescaling lowers the contribution of certain useful training samples, particularly for those with higher rewards. Moreover, deploying proximity optimization-based RL solutions requires additional resources to hold a reference model, which further increases the computational complexity. It is also worth noting that the aforementioned approaches mainly focus on mathematical problems or code generation, which are typical one-turn generation scenarios. When being adapted to multi-turn complex question answering, the algorithmic and implementation inefficiencies will become more pronounced.
In this work, we approach the design of a question answering agent from a conversational perspective, and devise our solution Mujica: a Multi-hop Joint Intelligence for Complex Question Answering. Mujica consists of two modules: 1) planner, and 2) worker, each of which serve as a user and an assistant for each other. The planner extracts subquestions, organizes them into a directed acyclic graph (DAG), and iteratively calls the worker to solve each subquestion based on their interdependencies. The worker serves as a mini-RAG system responsible for interacting with external databases and answering each subquestion assigned by the planner. This design allows conditionally independent subquestions to be resolved simultaneously, thereby improving efficiency.
Furthermore, we rethink the cons of policy gradient optimization in RL, and accordingly propose Minimalist policy Gradient Optimization (MyGO) to optimize Mujica. Unlike previous methods focusing on improving the policy gradient designs such as clipping importance sampling ratios schulman2017proximal or tweaking the advantage functions, MyGO shifts the burden to the improvement of sampling strategy. By sampling trajectories from the asymptotically approximate optimal policy, we can directly use Maximum Likelihood Estimation (MLE) for policy updates, which allows MyGO to be easily implemented with standard supervised finetuning frameworks. In addition, the utilization of MLE makes MyGO stable in training, and hyper-parameter-tuning-friendly. We provide detailed theoretical justification for the validity of MyGO. Our empirical experiments on various datasets show that Mujica-MyGO effectively improves the multi-hop question answering performance for different LLMs.
2 Question Answering Agent
2.1 Problem Definition
@@
This work addresses the task of Open-Domain Multi-Hop Question Answering (MHQA). QA systems aim to provide correct answers to given questions. Compared with single-hop QA, MHQA necessitates reasoning across multiple sources (hops) mavi2024multi , which renders MHQA a more challenging task. We focus on the open-domain setting, where relevant information must be retrieved from large-scale corpora (knowledge graphs or text collections). Details are provided in the Appendix.
@@
The inherent complexity of MHQA, characterized by multi-stage planning requirements and iterative evidence synthesis, necessitates framing the solution mechanism as an autonomous agent. We define an agent as an entity capable of planning, taking actions based on the environment, and engaging in sequential reasoning liu2025advances . We observe that LLMs equipped with RAG perform functions analogous to such an agent. Specifically, LLMs determine information requirements and search strategies (planning), formulate search queries (acting), and synthesize final answers from gathered evidence (reasoning). Recognizing this functional correspondence, we position the LLM itself as the core agent within our MHQA system.
Notation
Let $\mathcal{Q}$ be the set of all multi-hop questions, $\mathcal{Q}^{\prime}\subset\mathcal{Q}$ be the set of all sing-hop questions, $\mathcal{M}$ be the set of all LLM agents, $\mathcal{G}$ be the set of all dependency graphs. We define the answer of question $q\in\mathcal{Q}$ as a probability distribution $a_{q}$ and actions as functions associated with agents, e.g., $\mathtt{act}(\cdot|m)$, where $\mathtt{act}$ is a placeholder for actual actions and $m\in\mathcal{M}$. When the context is clear, we omit the notation of agent, i.e., we denote simply by $\mathtt{act}(\cdot)$.
2.2 Multi-hop Joint Intelligence for Complex Question Answering (Mujica)
@@
Figure 1 shows the overall pipeline of the proposed agentic QA framework. Given a complex question $q$, our QA agent aims to decompose it into a series of simple subquestions $\{q_{1},q_{2},\cdots,q_{n}\}$, and progressively answers each subquestion based on their internal dependencies until obtaining the final answer $a$. Let $\mathcal{I}$ denotes the instruction prompts, $\mathcal{C}=\{\mathcal{C}_{1},\mathcal{C}_{2},\cdots,\mathcal{C}_{n}\}$ denote the supporting contexts of each subquestion $q_{i}$, and $a_{i}$ an answer to it. Inspired by Search-o1 , we can formulate the answer generation process as the following objective function:
${}f_{\text{QA}}(q,I,\mathcal{C})=P(a|q,I,\mathcal{C})=P(q_{1}|q,\mathcal{I})\cdot\prod_{i=2}^{n}{P(q_{i}|a_{<i},q_{<i},q,\mathcal{I})}\cdot\prod_{i=1}^{n}{P(a_{i}|q_{i},\mathcal{C}_{i},\mathcal{I})},$ (1)
@@
where $q_{<i}$ and $a_{<i}$ denotes all preceding subquestions of subquestion $q_{i}$ and their corresponding answers. Based on the equation above, a QA agent can be naturally decoupled into two specialized modules that will be introduced in the following subsections:
@@
    Planner. Being responsible for planning how to answer a complex question, the planner determines the series of subquestions to be addressed, identifies what specific information needs to be retrieved, and adjusts the plan based on the answer of each subquestion.
@@
    Worker. Acting as a mini-RAG system, the worker directly interacts with the retriever and tackles specific sub-questions identified by the planner.
Planner and worker interact with each other through a conversational process, each serving as both user and assistant for one another. We denote our agentic QA framework as Multi-hop Joint Intelligence for Complex Question Answering (Mujica).
@@
Recent work (e.g., Search-o1 Search-o1 , PoG PoG ) has demonstrated that performing chain-of-thought reasoning through Eq. 1 is capable of deriving correct answers. However, in real-world scenarios, the subquestions posed are not naturally shaped as a sequence, since not all subquestions depend on answers to preceding ones. As shown in Fig. 2, there might be two conditionally independent subquestions, $\text{S}_{2,1}$ and $\text{S}_{2,2}$, which are dependent in the later subquestion $\text{S}_{4,1}$. The dependency relations form a directed acyclic graph (DAG). Directly applying Eq. 1 to such a reasoning process can be both inefficient and suboptimal. Therefore, to effectively handle DAG dependency graphs, we allow Mujica planner to ask subquestions in multiple iterations sequentially, where the answers of subquestions at an iterations can become dependent in later iterations. In each iteration, Mujica will simultaneously ask multiple conditionally independent subquestions.
Figure 2: QA as DAG.
@@
As aforementioned, the worker is responsible for handling interactions with the external environments, and answers questions assigned by the planner. Specifically, we utilize the same LLM-agent with a mini-RAG system. For each simple question $q_{i}$, the worker invokes an external retriever to fetch the top $k$ most relevant supporting contexts from a KG or a document corpus.  [^1] We employ bge-large-en-v1.5 sentence transformer as the external retriever. Then, the worker employs the LLM to review the target of the question, examine each retrieved context and accordingly anwser the subquestions outsourced by the planner. Specifically, the LLM is instructed to explicitly illustrate its thoughts, and select out relevant contexts by providing their indexes. Detailed prompts for planner and worker are provided in the Appendix.
3 Reinforcement Finetuning with Minimalist Policy Gradient Optimization
Figure 3: The proposed Minimalist Policy Gradient Optimization framework.
3.1 Preliminary: Reinforcement Learning Setting
Our QA agent is implemented within a conversational framework. A conversation (chat) is modeled as a finite sequence of interactions between the environment and the agent, denoted as $(e_{0},a_{0},e_{1},a_{1},\dots,e_{T},a_{T})$. The initial environment message $e_{0}$ comprises the user’s question and any necessary system prompts. For each turn $t\in[0,T]$, $a_{t}$ is agent’s action, and $e_{t+1}$ is the subsequent environment message. We define the state $s_{t}$ at turn $t$ as the history of all preceding interactions, $s_{t}=(e_{0},a_{0},e_{1},a_{1},\dots,a_{t-1},e_{t})$. Then, a trajectory $\tau=(s_{0},a_{0},s_{1},a_{1},\dots,s_{T},a_{T})$ to the final answer of the question can be denoted as the sequence of states and actions in each step. This formulation adheres to a Markov Decision Process (MDP). Since the golden reasoning path to the ground truth answer is not available in real-world scenarios and most datasets, we do not have supervision signals for each action $a_{t}$ to conduct supervised finetuning. The only supervision signal available to estimate the reward $r(\tau)\in\mathbb{R}$ for the entire trajectory $\tau$ is the final answer, which naturally corresponds to terminal reward in an RL setting.
We consider a policy-based method - the QA agent adopts a policy $\pi_{\theta}$ to generate reasoning trajectories, where the parameters $\theta$ are optimized using RL techniques. The objective of the RL is to find the best policy that maximizes the expected cumulative reward:

### v3

# Advancing Multi-Agent RAG Systems with Minimalist Reinforcement Learning

@@
Abstract.
Large Language Models (LLMs) equipped with modern Retrieval-Augmented Generation (RAG) systems often employ multi-turn interaction pipelines to interface with search engines for complex reasoning tasks. However, such multi-turn interactions inevitably produce long intermediate contexts, as context length grows exponentially with exploration depth. This leads to a well-known limitation of LLMs: their difficulty in effectively leveraging information from long contexts. This problem is further amplified in RAG systems that depend on in-context learning, where few-shot demonstrations must also be included in the prompt, compounding the context-length bottleneck. To address these challenges, we propose Mujica-MyGo, a unified framework for efficient multi-turn reasoning in RAG. Inspired by the divide-and-conquer principle, we introduce Mujica (Multi-hop Joint Intelligence for Complex Question Answering), a multi-agent RAG workflow that decomposes multi-turn interactions into cooperative sub-interactions, thereby mitigating long-context issues. To eliminate the dependency on in-context learning, we further develop MyGO (Minimalist Policy Gradient Optimization), a lightweight and efficient reinforcement learning algorithm that enables effective post-training of LLMs within complex RAG pipelines. We provide theoretical guarantees for MyGO’s convergence to the optimal policy. Empirical evaluations across diverse question-answering benchmarks—covering both text corpora and knowledge graphs—show that Mujica-MyGO achieves superior performance.
Key words and phrases:
Reinforcement Learning, Multi-Agent, RAG, QA, LLM
Equally contributed; work done during the internship at Huawei Noah’s Ark Lab.
1. Introduction
@@
The advent of Large Language Models (LLMs) ouyang2022training; achiam2023gpt has redefined expectations for Artificial Intelligence, offering new opportunities for many domains. Despite their versatility, LLMs suffer from hallucination huang2025survey and knowledge cutoffs cheng2024dated. While injecting knowledge through continuous learning is a possible solution, it is computationally expensive and risks catastrophic forgetting kirkpatrick2017overcoming. To address this issue, Retrieval-Augmented Generation (RAG) guu2020retrieval; lewis2020retrieval, a technique enabling LLMs to incorporate external knowledge via prompts, has drawn significant attention among researchers.
Most advanced RAG systems PoG; Search-o1; RoG; shao2023enhancing; khattab2022demonstrate employ a multi-turn pipeline to effectively address complex questions. This process typically involves decomposing the initial question into sub-questions, retrieving relevant information for each, and synthesizing a final answer. Such an approach allows for the generation of more fine-grained queries, enabling search engines to return more relevant passages and ultimately enhancing overall performance.
Despite its effectiveness, this multi-turn interaction introduces a significant challenge: the long-context problem. For instance, to answer the query, “Which U.S. presidents previously served as governors?”, a RAG system must first identify all U.S. presidents and then retrieve biographical information to determine which individuals held gubernatorial positions before their presidency. Each president has an extensive historical record, and aggregating such information results in a large volume of retrieved text within a single LLM context window. Given that LLMs struggle to capture information accurately in long contexts liu2024lost, this naive aggregation can lead to performance degradation.
Moreover, this issue is compounded by the reliance of In-Context Learning (ICL) ICL to guide the LLM’s behavior in most RAG workflows. This technique requires providing few-shot demonstrations within the prompt to instruct the model on the desired task. The performance of ICL is highly dependent on the precision and comprehensiveness of these examples, especially their ability to cover corner cases. Consequently, these detailed demonstrations further inflate the context length, exacerbating the long-context problem.
To address the aggregation problem, we propose Mujica (Multi-hop Joint Intelligence for Complex Question Answering), a multi-agent RAG system designed for complex queries. Mujica consists of two core modules: a Planner and a Worker. The Planner initiates the process by analyzing the original query, decomposing it into sub-questions, and delegating them to the Worker. Based on the responses from the Worker, the Planner iteratively formulates new queries until it has sufficient information to finalize an answer or reaches the iteration limit. The Worker, in turn, receives sub-questions from the Planner, retrieves relevant passages from search engines, and returns concise summaries. This planner-worker architecture enables a clear separation between high-level reasoning and low-level retrieval. By maintaining a clean, summary-based history in the Planner’s context, Mujica effectively mitigates the long-context problem.
To eliminate the need for cumbersome few-shot demonstrations in in-context learning (ICL), post-training is a crucial step for a complex RAG system like Mujica. However, the highly customized nature of such systems means that unified datasets for Supervised Fine-Tuning (SFT) are generally unavailable. This scarcity of trajectory data naturally frames the post-training process as a Reinforcement Learning (RL) problem, where the RAG system constitutes the policy and its performance on a task determines the reward. Nevertheless, optimizing this policy via RL is non-trivial due to the intricate interactions between the agents and the search engine. To tackle this challenge, we introduce Minimalist policy Gradient Optimization (MyGO), a simple and efficient RL algorithm specifically designed for such complex RAG systems.
Figure 1. The end-to-end architecture of the proposed Mujica framework.
We designed MyGO to be minimalist, deliberately avoiding components like reward weighting, importance sampling, or token clipping that are common in other algorithms. This approach stands in contrast to methods like Proximal Policy Optimization (PPO) schulman2017proximal, which requires an auxiliary value network to stabilize training, and Group Relative Policy Optimization (GRPO) shao2024deepseekmath, which must generate multiple trajectories per query for reward normalization, exacerbating latency. The key innovation in MyGO lies in its sampling strategy. By sampling trajectories from an asymptotically approximate optimal policy, we can directly use Maximum Likelihood Estimation (MLE) for policy updates. This insight allows MyGO to be implemented easily within standard SFT frameworks, lending it high stability during training and simplifying hyperparameter tuning. We provide a detailed theoretical justification for the validity of this approach and show through empirical experiments on various datasets that Mujica-MyGO effectively improves RAG performance across different LLMs.
2. RAG Agent
2.1. Problem Definition
@@
This work addresses Open-Domain Multi-Hop Question Answering (MHQA), a key challenge and benchmark for RAG. QA systems aim to provide correct answers to given questions. Compared with single-hop QA, MHQA necessitates reasoning across multiple sources (hops) mavi2024multi, which renders MHQA a more challenging task. We focus on the open-domain setting, where relevant information must be retrieved from large-scale corpora (knowledge graphs or text collections). Details are provided in Appendix A.
@@
The inherent complexity of MHQA, characterized by multi-stage planning requirements and iterative evidence synthesis, necessitates framing the solution mechanism as an autonomous agent. We define an agent as an entity capable of planning, taking actions based on the environment, and engaging in sequential reasoning liu2025advances. We observe that LLMs equipped with RAG perform functions analogous to such an agent. Specifically, LLMs determine information requirements and search strategies (planning), formulate search queries (acting), and synthesize final answers from gathered evidence (reasoning). Recognizing this functional correspondence, we position the LLM itself as the core agent within our system.
2.2. Mujica
As previously noted, repeated retrieval from RAG systems generates extended contexts that are challenging for LLMs to effectively process. To mitigate this, a straightforward approach is to adopt a divide-and-conquer approach by partitioning the context into smaller segments. Accordingly, we introduce Multi-hop Joint Intelligence for Complex Question Answering (Mujica) – a multi-agent RAG framework that decomposes multi-turn interactions into coordinated sub-interactions.
@@ 2.2. Mujica
Figure 1 shows the overall pipeline of the proposed agentic RAG framework. Given a complex question $q$, our RAG agent aims to decompose it into a series of simple subquestions $\{q_{1},q_{2},\cdots,q_{n}\}$, and progressively answers each subquestion based on their internal dependencies until obtaining the final answer $a$. Let $\mathcal{I}$ denote the instruction prompts, $\mathcal{C}=\{\mathcal{C}_{1},\mathcal{C}_{2},\cdots,\mathcal{C}_{n}\}$ denote the supporting contexts of each subquestion $q_{i}$, and $a_{i}$ an answer to it. Inspired by Search-o1, we can formulate the answer generation process as the following objective function:
$\penalty 10000\ \begin{aligned} &f_{\text{QA}}(q,I,\mathcal{C})=P(a|q,I,\mathcal{C})\\ &=\prod_{i=1}^{n}{P(q_{i}|a_{<i},q_{<i},q,\mathcal{I})}\cdot\prod_{i=1}^{n}{P(a_{i}|q_{i},\mathcal{C}_{i},\mathcal{I})},\end{aligned}$ (1)
@@
where $q_{<i}$ and $a_{<i}$ denote all preceding subquestions of subquestion $q_{i}$ and their corresponding answers. Based on the equation above, the QA framework can naturally be decoupled into two specialized roles:
@@
    Planner: Being responsible for planning how to answer a complex question, the planner determines the series of subquestions to be addressed, identifies what specific information needs to be retrieved, and adjusts the plan based on the answer of each subquestion.
@@
    Worker: Acting as a mini-RAG system, the worker directly interacts with the retriever and tackles specific sub-questions identified by the planner.
Planner and worker interact with each other through a conversational process, each serving as both user and assistant for one another. The planner and worker may operate as independent agents (distinct LLMs) or as virtual agents backed by a shared LLM but conditioned on distinct instruction prompts. Considering the substantial size of LLMs, we opt for the latter design for better computational efficiency.
@@
Recent work (e.g., Search-o1 Search-o1, PoG PoG) has demonstrated that performing chain-of-thought reasoning through Eq. 1 is capable of deriving correct answers. However, in real-world scenarios, the subquestions posed are not naturally shaped as a sequence, since not all subquestions depend on answers to preceding ones. As shown in Fig. 2, there might be two conditionally independent subquestions, $\text{S}_{2,1}$ and $\text{S}_{2,2}$, which are dependent in the later subquestion $\text{S}_{4,1}$. The dependency relations form a directed acyclic graph (DAG). Directly applying Eq. 1 to such a reasoning process can be both inefficient and suboptimal. Therefore, to effectively handle DAG dependency graphs, we allow our Mujica planner to ask subquestions in multiple iterations sequentially, where the answers of subquestions at an iteration can become dependent in later iterations. In each iteration, Mujica will simultaneously ask multiple conditionally independent subquestions.
@@
Figure 2. Modeling Complex Question Answering Process as a Directed Acyclic Graph.
@@
As aforementioned, the Worker is responsible for handling interactions with the external environments, and answers questions assigned by the planner. Specifically, we utilize the same LLM-agent with a mini-RAG system. For each simple question $q_{i}$, the worker invokes an external retriever to fetch the top $k$ most relevant supporting contexts from a KG or a document corpus.  [^1] We employ bge-large-en-v1.5 sentence transformer as the external retriever. Then, the Worker employs the LLM to review the target of the question, examine each retrieved context and accordingly anwser the subquestions outsourced by the Planner. Specifically, the LLM is instructed to explicitly illustrate its thoughts, and select relevant contexts by providing their indexes. Detailed prompts for Planner and Worker are provided in Appendix G.
3. RL for RAG Agent
Figure 3. The Proposed Minimalist Policy Gradient Optimization Framework.
3.1. Reinforcement Learning Setting
Our RAG agent is implemented within a conversational framework. A conversation is modeled as a finite sequence of interactions between the environment and the agent, denoted as $(e_{0},a_{0},e_{1},a_{1},\dots,e_{T},a_{T})$. The initial environment message $e_{0}$ comprises the user’s question and any necessary system prompts. For each turn $t\in[0,T]$, $a_{t}$ is the agent’s action, and $e_{t+1}$ is the subsequent environment message. We define the state $s_{t}$ at turn $t$ as the history of all preceding interactions, $s_{t}=(e_{0},a_{0},e_{1},a_{1},\dots,a_{t-1},e_{t})$. Then, a trajectory $\tau=(s_{0},a_{0},s_{1},a_{1},\dots,s_{T},a_{T})$ to the final answer of the question can be denoted as the sequence of states and actions in each step. This formulation adheres to a Markov Decision Process (MDP). Since the golden reasoning path to the ground truth answer is not available in real-world scenarios and most datasets, we do not have supervision signals for each action $a_{t}$ to conduct supervised finetuning. The only supervision signal available to estimate the reward $r(\tau)\in\mathbb{R}$ for the entire trajectory $\tau$ is the final answer, which naturally corresponds to the terminal reward in an RL setting.
We consider a policy-based method - the RAG agent adopts a policy $\pi_{\theta}$ to generate reasoning trajectories, where the parameters $\theta$ are optimized using RL techniques. The objective of the RL is to find the best policy that maximizes the expected cumulative reward: